Grade 128 grade maths

Calculate the amount and compound interest on
₹ 10800 for 3 years at 12𝟏 𝟐 % per annum compounded annually.

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Principal (P) = ₹ 10, 800
Rate (R) =12 whole 1/2 % =25/2 % (annual)
Number of years (n) = 3
Amount, A =P (1 +R/100 )^n
=₹[10800(1+25/200 )^3]
=₹ [10800 (225/200)^3]
=₹15377.34375
=₹15377.34 (approximately)
C.I. = A − P = ₹ (15377.34 − 10800) = ₹ 4,577.34

Thanks